# Bias Detection via Maximum Subgroup Discrepancy

### v1

@@ Abstract
Bias evaluation is fundamental to trustworthy AI, both in terms of checking data quality and in terms of checking the outputs of AI systems. In testing data quality, for example, one may study a distance of a given dataset, viewed as a distribution, to a given ground-truth reference dataset. However, classical metrics, such as the Total Variation and the Wasserstein distances, are known to have high sample complexities and, therefore, may fail to provide meaningful distinction in many practical scenarios.
@@
In this paper, we propose a new notion of distance, the Maximum Subgroup Discrepancy (MSD). In this metric, two distributions are close if, roughly, discrepancies are low for all feature subgroups. While the number of subgroups may be exponential, we show that the sample complexity is linear in the number of features, thus making it feasible for practical applications. Moreover, we provide a practical algorithm for the evaluation of the distance, based on Mixed-integer optimization (MIO). We also note that the proposed distance is easily interpretable, thus providing clearer paths to fixing the biases once they have been identified. It also provides guarantees for all subgroups. Finally, we empirically evaluate, compare with other metrics, and demonstrate the above properties of MSD on real-world datasets.
1 Introduction
@@
Regulatory frameworks, such as the AI Act [41] in Europe, suggest that one needs to measure data quality, including bias detection in training data, as well as to detect bias in the output of the AI system, but provide no suggestions as to what bias measures to use. This is the case of the very recent IEEE Standard for Algorithmic Bias Considerations [24] and earlier NIST white papers [42] too, where the latter stops at the “majority of fairness metrics are observational as they can be expressed using probability statements involving the available random variables”.
@@
At the most basic level, one could imagine bias detection as a two-sample problem in statistics, where, given two sets of samples, one asks whether they come from the same distribution. In practice, the two sets of samples often do not come from the same distribution, but one would like to have an estimate of the distance between the two distributions. The distance estimate, as any other statistical estimate [48], comes with an error. One would like the error in the estimate to be much smaller than the estimated value for the bias detection to be credible, stand up in any court proceedings, etc.
@@
In intersectional fairness, one would like to estimate bias for all protected subgroups. This faces three challenges: (1) the number of subgroups is exponential in the number of protected attributes, (2) estimating many distances on measure spaces has sample complexity exponential in the ambient dimension for each subgroup, and (3) some subgroups may have too few samples to estimate the bias well in terms of estimation error and significance. Let us now consider these challenges in more detail.
@@
Sample complexity is the number of samples that makes it possible to estimate a quantity to a given error. A lower bound on sample complexity then suggests the largest known number of samples universally required to reach a given error. The sample complexity of bias estimation depends on the distance between distributions (measures) used. The accuracy improves in the number of samples taken, but the rate of improvement depends on the dimension. See Table 1 for a brief summary of the commonly considered distances and their sample complexity and Section 2.1 for references and discussion. As is often the case in high-dimensional probability, the “curse of dimensionality” suggests that the number of samples to a given error grows exponentially with the dimension.
@@
Not only is sample complexity a lower bound on the runtime [33], but one must often test for bias for all subgroups, of which there are exponentially many (all combinations of protected attributes). Together, this could lead to doubly exponential time complexity, becoming intractable even for a few dimensions. Thus, one would like to detect bias for a single subgroup with polynomial complexity or reason about the joint problem. Additionally, the number of samples representing a certain subgroup decreases with the number of protected attributes defining it. Smaller sample complexity enables us to reliably detect bias for smaller subgroups.
@@
In this paper, we (i) introduce the Maximum Subgroup Discrepancy (MSD) distance between distributions. Compared to the standard distances discussed above, we (ii) show that this distance has a manageable sample complexity (linear in the number of protected attributes). We also note that in contrast to other distances, the MSD is interpretable, and thus provides actionable information on how the bias in the data should be mitigated once found. It also provides a guarantee in the form of an upper bound on bias over all subgroups. Next, we (iii) develop a new Mixed-integer optimization (MIO) formulation for the evaluation of MSD. Finally, we (iv) numerically validate the estimation stability and the dependence on the sample size for MSD and demonstrate its advantage over the Total Variation and Wasserstein metrics on 10 real-world datasets, which were developed with the view of studying fairness. We show that MSD requires exponentially fewer samples to obtain a good estimate of the true value of the distance. Taken together, we believe that these contributions suggest MSD as a theoretically grounded and practically useful bias detection method.
1.1 Problem description
@@
We now discuss the MSD metric in slightly more detail. Suppose we are given two probability distributions $\mu$ and $\nu$ over an input space $\mathcal{X}\subseteq\mathbb{R}^{d}$. A subset of the features, denoted $\mathcal{P}\subseteq[d]$, will be considered to be the protected attributes, such as for instance age, gender, ethnicity, and education. A subgroup is a subset of a population with a given fixed value of one or more protected attributes (e.g., black women). Formally, a subgroup $S$, defined by attributes $\mathcal{P}$ and their values $V=\left\{v_{p}\right\}_{p\in\mathcal{P}}\subseteq\mathbb{R}^{\left|\mathcal{P}\right|}$, is the set
$\mathcal{S}_{\mathcal{P}}=\left\{x=(x_{1},\ldots,x_{d})\in\mathcal{X}\hskip 5.69054pt|\hskip 5.69054ptx_{p}=v_{p}\text{ for $p\in\mathcal{P}$}\right\}.$ (1)
Two distributions will be considered similar if all subgroups have similar weights in the two distributions. Specifically, we define the Maximum Subgroup Discrepancy as
$\textrm{MSD}_{\Delta}(\mu,\nu;\mathcal{P})=\sup_{{S\in\mathcal{S}_{\mathcal{P}}}}\Delta({\mu(S),\nu(S)}),$ (2)
where $S$ ranges over all possible subgroups and $\Delta$ behaves similar to a distance on measure spaces over the input space, $\Delta:\mathcal{M}(\mathcal{X})\times\mathcal{M}(\mathcal{X})\to\mathbb{R}$. In this paper, we assume that the attributes are categorical [^1] This is always possible by applying standard quantization procedures. and thus the set of possible subgroups is finite. However, as mentioned earlier, this set will generally be of exponential size in the number of attributes.
@@
For example, in evaluating data quality, $\mu$ could be training data, seen as an empirical distribution and $\nu$ could be a distribution based on census data. Alternatively, we can consider distributions of a target variable of an ML model in auditing its fairness: $\mu$ for the positive class and $\nu$ for the negative. We can then find a subgroup that is most disadvantaged (in terms of $\textrm{MSD}_{\Delta}(\mu,\nu;\mathcal{P})$) by the model.
@@
2 Background and Related work
@@
Table 1: Sample complexity of estimating popular distances on measure spaces in terms of the ambient dimension $d$.
Distance Samples Ref.
Total Variation $\mathrm{\Omega}(\exp(d))$ [11, 54]
Hellinger (Jeffreys) $\mathrm{\Omega}(\exp(d))$ [11, 54]
Wasserstein-1 $\infty$ Thm 5 [33]
Wasserstein-2 $\mathrm{\Omega}(\exp(d))$ [13, 17, 52]
@@
Wasserstein-$\infty$ $\mathrm{\Omega}(\exp(d))$ [17, 34, 52]
Maximum Mean Discrepancy $\mathrm{\Omega}(\exp(d))$ [47]
Maximum Subgroup Discrepancy $\mathrm{O}(d)$ w.h.p. This paper (Eq. 13 in Section 3.1)
2.1 Distances on Measure Spaces
@@
There are numerous distances on measure spaces used in applied probability, including (in the approximately chronological order) Total Variation (TV, [25]), Hellinger distance [22], Kullback–Leibler (KL) divergence [31] and its variants, Wasserstein-2 [50, 13] and its variants such as Wasserstein-1 [50], and Maximum Mean Discrepancy (MMD, [19]). As we outline below, most of these methods have exponential sample complexity in terms of the number of dimensions. As can be seen in Table 1, most of these distances come with either undecidability results [33] or exponential lower bounds on their sample complexity in the worst case. For MMD, while [19] claimed polynomial sample complexity, [47] explained the lower bounds on sample complexity under strong assumptions. Yet for other methods, such as Hellinger and Jeffreys distances, high sample complexity follows from their relationship to TV distance. We refer to [40] for a thorough survey.
@@
Obviously, one can consider additional assumptions, such as having cardinality of the support (which scales with $\exp(d)$ in general), bounded by a constant. Testing TV closeness in time sublinear in the support was derived in [4]. Entropy estimation bounds were obtained in [49]. [14, 2] give an algorithm for estimating TV between just product measures, polynomial in product dimension. Likewise, one can consider smoothness of the measures and certain invariance properties [5, 46], or focus on Gaussian distributions [23] only. Ising type models testing was explored in [26]. While these assumptions are of considerable interest, it is not easy to test that those assumptions hold in real-world data sets.
2.2 Subgroup and Intersectional Fairness
The notion of subgroups gave rise to the work on subgroup fairness [27], and underlies the work on intersectional fairness [16, 18]. In particular, in the legal scholarship, Intersectional Fairness ideas go back to the work of Crenshaw [8], but remain a subject of lively debate [7] to the present day.
@@ 2.2 Subgroup and Intersectional Fairness
In the algorithmic fairness literature, the sample complexity of certain fairness estimates (statistical parity, false positive fairness) was considered in [27]. The notion of distance, which we consider here, is conceptually different, as it concerns data quality (see Section 1) rather than a fairness test of a particular given classifier. Moreover, the algorithms of [27] were developed with linear subgroups in mind, and consequently evaluated on linear subgroups [28]. In particular, these algorithms require certain specific heuristics (oracles), that are mainly developed for the linear case. We note that such linear subgroups are considerably less interpretable and less suitable for real-world applications compared to the intersectional subgroups that we consider here.
@@
In a related direction, multidimensional subset scanning methods systematically sift through potentially large collections of subgroups to find anomalous or biased regions. Although these approaches vary in the exact objective – ranging from pinpointing classifier miscalibration [56] to detecting compact anomalies via penalties [44] or scanning multiple data streams [39] – they share a focus on likelihood-based scoring and efficient “fast subset” searches. While these techniques are adept at finding one or more “highest-scoring” subgroups, their goals and scoring mechanisms are distinct from distance-based comparisons, positioning them as complementary.
2.3 Learning DNF
@@
The study of learning DNF formulas is a fundamental part of theoretical computer science. Despite its polynomial sample complexity, finding a general algorithm with polynomial-time complexity has proven elusive in many settings. Since the breakthrough reduction to polynomial threshold functions of [29], which matched the 1968 lower bound of [36], [9] helped understand the complexity-theoretic barriers. See also [43] for an overview.
2.4 Mixed-integer optimization

### v2

@@ Abstract.
Bias evaluation is fundamental to trustworthy AI, both in terms of checking data quality and in terms of checking the outputs of AI systems. In testing data quality, for example, one may study the distance of a given dataset, viewed as a distribution, to a given ground-truth reference dataset. However, classical metrics, such as the Total Variation and the Wasserstein distances, are known to have high sample complexities and, therefore, may fail to provide a meaningful distinction in many practical scenarios.
@@
In this paper, we propose a new notion of distance, the Maximum Subgroup Discrepancy (MSD). In this metric, two distributions are close if, roughly, discrepancies are low for all feature subgroups. While the number of subgroups may be exponential, we show that the sample complexity is linear in the number of features, thus making it feasible for practical applications. Moreover, we provide a practical algorithm for evaluating the distance based on Mixed-integer optimization (MIO). We also note that the proposed distance is easily interpretable, thus providing clearer paths to fixing the biases once they have been identified. Finally, we describe a natural general bias detection framework, termed MSDD distances, and show that MSD aligns well with this framework. We empirically evaluate MSD by comparing it with other metrics and by demonstrating the above properties of MSD on real-world datasets.
intersectional fairness; Maximum Subgroup Discrepancy; DNF learning
journalyear: 2025conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736857copyright: ccisbn: 979-8-4007-1454-2/2025/08ccs: Theory of computation Complexity theory and logicccs: Mathematics of computing Multivariate statistics
1. Introduction
@@
Regulatory frameworks, such as the AI Act (Parliament and Council, 2024) in Europe, suggest that one needs to measure data quality, including bias detection in training data, as well as to detect bias in the output of the AI system, but provide no suggestions as to what bias measures to use. This is the case of the very recent IEEE Standard for Algorithmic Bias Considerations (IEEE Standards Association, 2025) and earlier NIST white papers (Schwartz et al., 2022) too, where the latter stops at the “majority of fairness metrics are observational as they can be expressed using probability statements involving the available random variables”.
@@
At the most basic level, one could imagine bias detection as a two-sample problem in statistics, where, given two sets of samples, one asks whether they come from the same distribution. In practice, the two sets of samples often do not come from the same distribution, but one would like to have an estimate of the distance between the two distributions. The distance estimate, as any other statistical estimate (Tsybakov, 2009), comes with an error. One would like the error in the estimate to be much smaller than the estimated value for the bias detection to be credible, stand up in any court proceedings, etc.
Figure 1. An illustration of intersectional bias. A classifier accepts an equal proportion of persons of each race and age group. Thus, demographic parity is achieved, and no such bias is detected. However, when considering subgroups, i.e., intersections of protected attributes, we see that young blue people are severely disadvantaged. This disparity is found by our proposed measure, Maximum Subgroup Discrepancy (MSD), which also recovers the description of the most disadvantaged subgroup (race = "Blue" AND age $\in$ [0, 18]).
A figure illustrating the importance of intersectional fairness. In two dimensions, age and race are binned into 3 by 5 bins, each having the same number of samples. Each sample is either classified as positive or negative. In each row and column is the same proportion of positives and negatives, which means that demographic parity is achieved. However, looking at the bin of young blue people, one notices that all of them are classified as negative. The proposed measure, MSD, is able to capture that and recover the most disadvantaged subgroup.
@@
In intersectional fairness, one would like to estimate bias for all protected subgroups. See Figure 1 for a motivating toy example. Estimation of intersectional bias faces three challenges: (1) the number of subgroups is exponential in the number of protected attributes, (2) estimating many distances on measure spaces has sample complexity exponential in the ambient dimension for each subgroup, and (3) some subgroups may have too few samples to estimate the bias well in terms of estimation error and significance. Let us now consider these challenges in more detail.
@@
Sample complexity is the number of samples that makes it possible to estimate a quantity to a given error. A lower bound on sample complexity then suggests the largest known number of samples universally required to reach a given error. The sample complexity of bias estimation depends on the distance between distributions (measures) used. The accuracy improves with the number of samples taken, but the rate of improvement depends on the dimension. See Table 1 for a brief summary of the commonly considered distances and their sample complexity, and Section 2.1 for references and discussion. As is often the case in high-dimensional probability, the “curse of dimensionality” suggests that the number of samples to a given error grows exponentially with the dimension.
@@
Not only is sample complexity a lower bound on the runtime (Lee et al., 2024), but one must often test for bias for all subgroups, of which there are exponentially many in the number of protected attributes. Together, this could lead to doubly exponential time complexity, challenging even for low dimensions. Thus, one would like to detect bias for a single subgroup with polynomial complexity or reason about the joint problem. Additionally, the number of samples representing a certain subgroup decreases with the number of protected attributes defining it. Smaller sample complexity enables us to reliably detect bias for smaller subgroups.
@@
In this paper, we (i) introduce the Maximum Subgroup Discrepancy (MSD) metric. Compared to the standard distances discussed above, we (ii) show that it has a manageable sample complexity (linear in the number of protected attributes). We also note that in contrast to other distances, the MSD is interpretable and thus provides actionable information on how the bias in the data should be mitigated once found. It also provides a guarantee in the form of an upper bound on bias over all subgroups. Next, we (iii) develop a new mixed-integer optimization (MIO) formulation for the evaluation of $\mathrm{MSD}$. Finally, we (iv) numerically validate the estimation stability and the dependence on the sample size for $\mathrm{MSD}$ and demonstrate its advantage over common distribution distance measures on 10 real-world datasets, which were developed with the view of studying fairness. We show that $\mathrm{MSD}$ requires exponentially fewer samples to obtain a good estimate of the true value of the distance. Taken together, we believe that these contributions suggest MSD as a theoretically grounded and practically useful bias detection method.
1.1. Problem description
@@
To arrive at the definition of MSD, we first discuss the most general intersectional bias problem. Suppose we are given two probability distributions $\mu$ and $\nu$ over an input space $\mathcal{X}\subseteq\mathbb{R}^{d}$. A subset of the features, denoted $\mathcal{P}\subseteq[d]$, will be considered to be the protected attributes, such as for instance age, gender, ethnicity, and education. A subgroup is a subset of a population with a given fixed value of one or more protected attributes (e.g., black women). Formally, a subgroup $S$, defined by attributes $\mathcal{P}$ and their values $V=\left\{v_{p}\right\}_{p\in\mathcal{P}}\subseteq\mathbb{R}^{\left|\mathcal{P}\right|}$, is the set
(1) $S=S_{\mathcal{P},V}=\left\{x=(x_{1},\ldots,x_{d})\in\mathcal{X}\hskip 5.69054pt|\hskip 5.69054ptx_{p}=v_{p}\text{ for $p\in\mathcal{P}$}\right\}.$
Denote by $\mathcal{S}_{\mathcal{P}}$ the set of all possible subgroups $S$. To study intersectional fairness, one would like to estimate the distance between the two distributions restricted to a given subgroup and find the subgroup for which these distributions differ most. Specifically, we call the Maximum Subgroup Distribution Distance
(2) $\mathrm{MSDD}_{\Delta}(\mu,\nu;\mathcal{P})=\sup_{{S\in\mathcal{S}_{\mathcal{P}}}}\Delta({\mu_{|S},\nu_{|S}}),$
where $\mu_{|S}$ defines the restriction measure to the subset $S$: $\mu_{|S}(A):=\mu(A\cap S)/\mu(S)$, and $\Delta$ a (possibly semi-) metric on measure spaces over the input space, $\Delta:\mathcal{M}(\mathcal{X})\times\mathcal{M}(\mathcal{X})\to\mathbb{R}$. For instance, $\Delta$ could be the Total Variation or the Wasserstein distances (Section 2.1).
@@
To give an example for the use of MSDD in evaluating data quality, $\mu$ could be training data, seen as an empirical distribution, and $\nu$ could be a distribution based on census data. Alternatively, we can consider distributions of a target variable of an ML model in auditing its fairness: $\mu$ for the positive class and $\nu$ for the negative. We can then find a subgroup that is most disadvantaged (in terms of $\mathrm{MSDD}_{\Delta}(\mu,\nu;\mathcal{P})$) by the model.
The MSDD is a natural general metric that may be used to quantify intersectional bias in a variety of scenarios, including both data quality and system output settings. However, it also has a number of fundamental practical issues: (i) The enumeration of all subgroups would have exponential time complexity in the number of protected attributes, and (ii) common distances on measure spaces ($\Delta$) have high sample complexity (Section 2.1), which makes them impossible to estimate for small subgroups. Note, however, that since the number of subgroups is large, some of the subgroups will almost inevitably be small. (iii) The run time complexity of many standard distribution metrics is superlinear (Nguyen and Ho, 2022; Gretton et al., 2012) which is then compounded by the need to compute the distances for many subgroups.
To alleviate these issues, we introduce the MSD distance. The MSD can be used as the $\Delta$ for the $\mathrm{MSDD}$ distance and is itself based on the notion of subgroups (the attributes $\mathcal{P}$ along which $\mathrm{MSDD}$ is computed would normally be disjoint from the ones used to define MSD). As discussed earlier, MSD has favorable sample complexity properties and an efficient computation algorithm, thus providing considerable stability and speed gains for issues (ii) and (iii).
@@
2. Background and Related work
@@
Table 1. Sample complexity of estimating popular distances on measure spaces in terms of the ambient dimension $d$.
Distance Samples Reference
Wasserstein-1 $\infty$ Thm 5 (Lee et al., 2024)
Total Variation $\mathrm{\Omega}(\exp(d))$ (Devroye et al., 2018; Wolfer and Kontorovich, 2021)
Hellinger (Jeffreys) $\mathrm{\Omega}(\exp(d))$ (Devroye et al., 2018; Wolfer and Kontorovich, 2021)
Wasserstein-2 $\mathrm{\Omega}(\exp(d))$ (Dudley, 1969; Fournier and Guillin, 2015; Weed and Bach, 2019)
@@
Wasserstein-$\infty$ $\mathrm{\Omega}(\exp(d))$ (Fournier and Guillin, 2015; Liu et al., 2019; Weed and Bach, 2019)
MMD $\mathrm{\Omega}(\exp(d))$ (Tolstikhin et al., 2016)
$\mathrm{MSD}$ $\mathrm{O}(d)$ w.h.p. This work (Eq. 13 in Sec. 3.1)
2.1. Distances on Measure Spaces
@@
There are numerous distances on measure spaces used in applied probability, including (in the approximately chronological order) Total Variation (TV, (Jordan, 1881)), Hellinger distance (Hellinger, 1909), Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) and its variants, Wasserstein-2 (Vaserstein, 1969; Dudley, 1969) and its variants such as Wasserstein-1 (Vaserstein, 1969), and Maximum Mean Discrepancy (MMD, (Gretton et al., 2012)). As we outline below, most of these methods have exponential sample complexity in terms of the number of dimensions. As can be seen in Table 1, most of these distances come with either undecidability results (Lee et al., 2024) or exponential lower bounds on their sample complexity in the worst case. For MMD, while (Gretton et al., 2012) claimed polynomial sample complexity, (Tolstikhin et al., 2016) explained the lower bounds on sample complexity under strong assumptions. Yet for other methods, such as Hellinger and Jeffreys distances, high sample complexity follows from their relationship to TV distance. We refer to (Panaretos and Zemel, 2019) for a thorough survey.
@@
Obviously, one can consider additional assumptions, such as having cardinality of the support (which scales with $\exp(d)$ in general), bounded by a constant. Testing TV closeness in time sublinear in the support was derived in (Chan et al., 2014). Entropy estimation bounds were obtained in (Valiant and Valiant, 2011). (Feng et al., 2023; Bhattacharyya et al., 2023) give an algorithm for estimating TV between just product measures, polynomial in product dimension. Likewise, one can consider smoothness of the measures and certain invariance properties (Chen et al., 2023; Tahmasebi and Jegelka, 2023), or focus on Gaussian distributions (Hsu et al., 2024) only. Ising-type models testing was explored in (Kandiros et al., 2023). While these assumptions are of considerable interest, it is not easy to test whether those assumptions hold in real-world data sets.
2.2. Subgroup and Intersectional Fairness
@@
The notion of subgroups gave rise to the work on subgroup fairness (Kearns et al., 2018), and underlies the work on intersectional fairness (Foulds et al., 2020; Gohar and Cheng, 2023). In particular, in the legal scholarship, Intersectional Fairness ideas go back to the work of Crenshaw (Crenshaw, 2013), but remain a subject of lively debate (Collins and Bilge, 2020) to the present day. In the algorithmic fairness literature, the sample complexity of certain fairness estimates (statistical parity, false positive fairness) was considered in (Kearns et al., 2018). The notion of distance, which we consider here, is conceptually different, as it concerns data quality (see Section 1) rather than a fairness test of a particular given classifier. Moreover, the algorithms of (Kearns et al., 2018) were developed with linear subgroups in mind, and consequently evaluated on linear subgroups (Kearns et al., 2019). In particular, these algorithms require certain specific heuristics (oracles) that are mainly developed for the linear case. We note that such linear subgroups are considerably less interpretable and less suitable for real-world applications compared to the intersectional subgroups that we consider here. Note that naively extending the definitions used in marginal fairness (considering only groups defined by a single protected attribute at a time (Dominguez-Catena et al., 2024)) to intersectional fairness will suffer from low sample sizes of small subgroups. Estimates using such extensions (Chen et al., 2024) will have dubious significance and will be less robust to outliers.
@@
In a related direction, multidimensional subset scanning methods systematically sift through potentially large collections of subgroups to find anomalous or biased regions. Although these approaches vary in the exact objective – ranging from pinpointing classifier miscalibration (Zhang and Neill, 2016) to detecting compact anomalies via penalties (Speakman et al., 2015) or scanning multiple data streams (Neill et al., 2013) – they share a focus on likelihood-based scoring and efficient “fast subset” searches. While these techniques are adept at finding one or more “highest-scoring” subgroups, their goals and scoring mechanisms are distinct from distance-based comparisons.
Finally, there is vast literature on the topic of debiased learning (e.g., Bahng et al., 2020; Nam et al., 2020; Ahn et al., 2023), which is concerned with learning unbiased models on biased data. Said bias is, however, in the form of spurious correlations, making the task distinct from the one considered here.
2.3. Learning DNF
@@
The study of learning DNF formulas is a fundamental part of theoretical computer science. Despite its polynomial sample complexity, finding a general algorithm with polynomial-time complexity has proven elusive in many settings. Since the breakthrough reduction to polynomial threshold functions of (Klivans and Servedio, 2001), which matched the 1968 lower bound of (Minsky and Papert, 1988), (Daniely and Shalev-Shwartz, 2016) helped understand the complexity-theoretic barriers. See also (Shalev-Shwartz and Ben-David, 2014) for an overview.
2.4. Mixed-integer optimization